OpenStudy (anonymous):

2x3 - 5x2y + xy2 + 2y3 a. Simplify (2x2 + y2)(x - 2y) = ? b. Simplify (2x + y)(x2 -3xy + 2y 2 ) = ?

OpenStudy (anonymous):

where is c and d?

OpenStudy (anonymous):

\[2\,{x}^{3}-5\,{x}^{2}y+x{y}^{2}+2\,{y}^{3}\] did you mean that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ooh where are c and d then :) i got my own answer can i see c and d?

OpenStudy (anonymous):

theyre not right for this

OpenStudy (anonymous):

so far i think it is b but i feel it can be simplified further :)

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

was it a?

OpenStudy (anonymous):

well i get :) (-y+x)*(x-2*y)*(y+2*x) fully factored

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

:) i'll check a and b :)

OpenStudy (anonymous):

i dont think it is a :(

OpenStudy (anonymous):

now for b

OpenStudy (anonymous):

its b!!!!!!!!!!!!!!!!!!:)))))))))))))))))))))))))))

OpenStudy (anonymous):

told ya ;)

OpenStudy (anonymous):

and i check to and it is b :D

OpenStudy (anonymous):

If you know algrabeic long division then there is a theory of if you divide your given polynomial in this case 2x3 - 5x2y + xy2 + 2y3 by a suspected binomial or trinomial (2x + y) or (x2 -3xy + 2y 2 ) ^ those are from B the 2 expressions inside the ( ) and you get 0 then that trinomial or binomial is a factor and therfore part of the answer :)

OpenStudy (anonymous):

if you get a remainder of 0 :P

OpenStudy (anonymous):

in other words 2x3 - 5x2y + xy2 + 2y3 -------------------- = (x2 -3xy + 2y 2 ) with no remainder (2x + y) and obviously 2x3 - 5x2y + xy2 + 2y3 --------------------- = 2x+y with no remainder (x2 -3xy + 2y 2 )

OpenStudy (anonymous):

But if you dont know algabreic long division dont bother with that method :P